OpenStudy (faman39):

Find the equivalent ratios. 9/5 21 to 35 27: 30 8/32

OpenStudy (anonymous):

none of them are equivalent

OpenStudy (faman39):

those are down, not this one

OpenStudy (faman39):

sorry i need help i forgot

OpenStudy (faman39):

in this

OpenStudy (across):

With time, this should come intuitively. However, I believe a good way to 'visualize' this type of problem is to decompose each ratio into its prime factors and simplify:\[\frac{9}{5}=\frac{3\cdot3}{5}\]\[\frac{21}{35}=\frac{3\cdot7}{5\cdot7}=\frac{3\cdot\not{7}}{5\cdot\not{7}}=\frac{3}{5}\]\[\frac{27}{30}=\frac{3\cdot3\cdot3}{2\cdot3\cdot5}=\frac{3\cdot\not{3}\cdot3}{2\cdot\not{3}\cdot5}=\frac{9}{10}\]\[\frac{8}{32}=\frac{2\cdot2\cdot2}{2\cdot2\cdot2\cdot2\cdot2}=\frac{\not{2}\cdot\not{2}\cdot\not{2}}{\not{2}\cdot\not{2}\cdot\not{2}\cdot2\cdot2}=\frac{1}{4}\]You can see that none of them are equivalent.

OpenStudy (anonymous):

\[\frac{9}{5}=\frac{21}{35}?\] \[35\times 9=5\times 51?\] \[175=225?\] NO

OpenStudy (anonymous):

\[\frac{9}{5}=\frac{27}{30}?\] \[\frac{9}{5}=\frac{9}{10}?\] \[9\times 10=5\times 9?\hell no

OpenStudy (anonymous):

in fact \[\frac{9}{5}>1\] but \[\frac{27}{30}<1\]

OpenStudy (anonymous):

across is right boss

Directrix (directrix):

NONE 9/5 = 1.8 21 to 35 = .6 27: 30 = .9 8/32 = .25